### v1

@@ IV Klein-Gordon equation
During 1926, a possible relativistic wave equation was derived by at least seven authors: Klein [52], Schrödinger [36], Fock [53], Donder, Dungen [54], Kudar [55] and Gordon [56]. Priority in the nomenclature is given to Oscar Klein, who proposed the equation in April 1926, while Fock provided a more intricate exploration of relativistic wave mechanics [3]. We will follow the more direct derivation, starting from the relativistic relation between energy $E$ and linear momentum $\boldsymbol{p}=(p_{1},p_{2},p_{3})$, expressed as
$E^{2}=(\boldsymbol{p}c)^{2}+(mc^{2})^{2},$ (36)
where the potential energy $V(\mathbf{r},t)$ is taken as zero and $m$ refers to the rest mass. We can replace both $E$ and $\boldsymbol{p}$ quantities in Eq. (36) by operators the operator in Eqs. (6) and (7), so that we obtain
@@
$\frac{1}{c^{2}}\frac{\partial^{2}\psi\left(\mathbf{r},t\right)}{\partial t^{2}}=\boldsymbol{\nabla}^{2}\psi\left(\mathbf{r},t\right)-\left(\frac{mc}{\hbar}\right)^{2}\psi\left(\mathbf{r},t\right),$ (38)
@@
Another crucial aspect of the KG equation is investigating its applicability in physical systems, particularly considering the range of problems that can be addressed with the Schrödinger equation. Therefore, applying the KG equation to a simple system and testing its validity is only fair. For instance, as an elementary problem, let us analyze the energy eigenvalues of a free particle. Taking Eq. (39) and simplifying it, we have
$\frac{1}{c^{2}}\frac{\partial^{2}\psi\left(\mathbf{r},t\right)}{\partial t^{2}}-\boldsymbol{\nabla}^{2}\psi\left(\mathbf{r},t\right)+\mu{{}^{2}}\psi\left(\mathbf{r},t\right)=0,$ (45)
which is expected to have plane-wave solutions with a time dependency of the type $-(i\varepsilon t)/\hbar$ and a spatial dependence of the type $+(i\boldsymbol{p}\cdot\mathbf{r})/\hbar$, so that
$\psi\left(\mathbf{r},t\right)=N\exp[-\frac{i}{\hbar}(\varepsilon t-\boldsymbol{p}\cdot\mathbf{r})],$ (46)
@@
$\varepsilon^{2}=\boldsymbol{p}^{2}c^{2}+\mu^{2}c^{2}\hbar^{2},$ (47)
@@
where we note that, although it is a preserved quantity, $\rho(\mathbf{r},t)$ is not positive definite, unlike Eq. (14). This posed an immediate challenge as it complicated the probabilistic interpretation of the wave function. Nevertheless, we now know that the correct interpretation is that $\rho(\mathbf{r},t)$ is a probability charge density [42, 482].
@@ V Pauli equation
An important event in the search for a relativistic wave equation was the discovery of spin, initially believed to be intrinsically relativistic. The first indication of spin’s existence arose from the Stern-Gerlach (SG) experiment [60], where an atomic beam of silver was subjected to an inhomogeneous magnetic field. Silver’s 47th electron’s magnetic moment approximates the atom’s magnetic moment because of its specific structure. If the electron behaved classically, its magnetic moment would exhibit a continuous range of values, resulting in a Gaussian pattern after passing through the magnetic field. However, in the SG experiment, deviations were observed for all atoms, with detection occurring in two preferential regions, as illustrated in Fig. 4. This suggested that the magnetic moment of the electron had only two directions, a phenomenon which Sommerfeld interpreted as evidence of space quantization [24, 131].
Figure 4: SG experiment illustrated schematically. The beam of silver atoms is expelled from the furnace, passes through a collimator, and is subjected to a heterogeneous magnetic field $\mathbf{B}$. In the detector, two preferential regions are perceived. Adapted from Ref. [42, 2].
@@
First, let us discuss how the Schrödinger equation deals with electromagnetic interactions by writing it as a function of $\boldsymbol{p}$ as
@@
$\boldsymbol{\Pi}=\boldsymbol{p}-\frac{e}{c}\mathbf{A},$ (50)
@@
$\boldsymbol{p}$ $\rightarrow\boldsymbol{\Pi}=\boldsymbol{p}-\frac{e}{c}\mathbf{A},$ (51)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left[\frac{1}{2m}\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)^{2}+e\phi\right]\psi\left(\mathbf{r},t\right).$ (53)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left\{\frac{1}{2m}\left[\mathbf{\boldsymbol{\sigma}}\ \cdot\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)\right]^{2}+e\phi\right\}\psi\left(\mathbf{r},t\right),$ (54)
@@
understood as operators associated with spin. In this equation, $\psi\left(\mathbf{r},t\right)$ is a two-component wave function [64, 340] – a mathematical entity called a spinor represented by a column matrix with two components. Incorporating the spinor into the wave function is a direct consequence of the binary nature of spin. Consequently, the solutions of the Pauli equation involve two differential equations, as opposed to one in the case of the Schrödinger equation. In this work, we will not go into the formal aspects of the spinor, which can be found in [65].
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left\{\frac{1}{2m}\left[\left(\boldsymbol{p}-\frac{e}{c}\mathbf{A}\right)^{2}-\frac{e\hbar}{c}\mathbf{\boldsymbol{\sigma}}\cdot\mathbf{B}\right]+e\phi\right\}\psi\left(\mathbf{r},t\right),$ (57)
where $\mathbf{B}=\boldsymbol{\nabla}\times\mathbf{A}$ is the magnetic field. In the context of our discussion, the significance of the Pauli equation must be underscored. By incorporating spin, even if forcefully, into the Schrödinger wave function, his model became the most accurate representation of an electron interacting with an electromagnetic field under a non-relativistic regime.
@@ VI Dirac equation
Dirac, dissatisfied with the KG equation, believed that a linear relativistic wave equation could be developed in terms of the temporal derivative. This conviction arose from transformation theory [16, 288], a concept to which he and Jordan contributed, aimed at generalizing both matrix and wave mechanics [24, 307]. Upon comparing the Schrödinger equation with the KG equation, it is apparent that the Schrödinger equation fulfills this criterion, whereas the KG equation does not. To understanding why this linearity is vital in the development of quantum theory, see [68, 69, 233]. Dirac expressed his strong preference for transformation theory with his statement: “The transformation theory had become my darling. I was not interested in considering any theory that would not fit my darling” [70] as cited in [16, 289]. Notably, since the Dirac equation is of utmost importance for this article, the steps concerning it will be elaborated upon in greater detail, with a deeper exploration of specific topics compared to the earlier models. The derivation presented here emphasizes Dirac’s initial insight, while a more didactic derivation can be found in [71] and more explicitly detailed in [13].
Thus, while “playing with the equations” [16, 290], Dirac obtained
$\mathbf{\boldsymbol{p}}^{2}\cdot\mathbbm{1}_{2}=(\boldsymbol{\sigma}\cdot\boldsymbol{p})^{2}.$ (58)
@@
Dirac claimed that both he and Pauli derived those matrices independently [16, 290].
@@
$\gamma_{\mu}\gamma_{\nu}={}$ $-\gamma_{\nu}\gamma_{\mu},$ for $\mbox{for }\mu\neq\nu.$
@@
which we can rewrite using the scalar product between $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $\boldsymbol{p}$ as
$E=c\boldsymbol{\alpha}\cdot\boldsymbol{p}+\beta mc^{2},$ (73)
@@
$i\hbar\frac{\partial}{\partial t}=c\boldsymbol{\alpha}\cdot\boldsymbol{p}+\beta mc^{2}.$ (74)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left(c\boldsymbol{\alpha}\cdot\boldsymbol{p}+\beta mc^{2}\right)\psi\left(\mathbf{r},t\right).$ (75)
@@
$H=c\boldsymbol{\alpha}\cdot\boldsymbol{p}+\beta mc^{2},$ (77)
@@
$i\hbar\frac{\partial\psi\left(\mathbf{r},t\right)}{\partial t}=\left(-i\hbar c\boldsymbol{\alpha}\cdot\nabla+\beta mc{{}^{2}}\right)\psi(\mathbf{r},t).$ (78)
@@
Just as we did with the KG equation, let us analyze the case of free particles. Firstly, we consider a particle at rest. Although it is a restricted case, it will provide us with the necessary insight regarding the allowed energy values. We employ the Dirac equation with $\boldsymbol{p}=0$, so that
@@
The first two expressions align with the anticipated outcomes for a free-particle scenario, as suggested by the negative sign in the exponent. However, as we will explore in the subsequent case, the two lower components deviate from the expected behavior and are associated with negative energy.
@@
After some confusion [16, 348], Dirac managed to formulate a coherent interpretation of this problem using the Pauli exclusion principle [74]. He conjectured a sea of negative energy, referred to as Dirac Sea, in which all the negative-energy states are filled, thus preventing an electron transition to negative energy. According to Dirac’s idea, a “hole” in this sea could behave like a particle with positive charge. Initially, Dirac speculated that the particle associated with the "hole" was the proton. This belief stemmed from the prevailing notion at the time that there were only two elementary particles in nature. However, as pointed out by Oppenheimer [75], the presence of a proton in this sea of negative energy would lead to annihilation and the release of energy in the form of photons, which means, in a practical sense, the instability of matter. Weyl [76], employing the symmetry in the Maxwell and Dirac equations, pointed out that the mass of the hole would have to be the same as the electron. In response to these critiques, Dirac stated, in an article published in 1931 [77], that if there is a hole, it represents a new, experimentally unknown particle with the same mass of the electron but opposite charge – an “antielectron”, so to speak. Today, this particle has been experimentally confirmed and is known as the positron. Therefore, the new degree of freedom observed in the bispinor arises precisely from the antiparticle associated with the particle. The energy diagram in Fig. 5 illustrates the concept of the Dirac Sea.
@@
Figure 5: Dirac equation’s energy diagram. Adapted from Ref. [42, 495].
@@
$\psi_{R}^{+}={}$ $u_{R}^{+}\exp\left[-\frac{i}{\hbar}\left(\varepsilon_{p}t-pz\right)\right],$ $\psi_{L}^{+}={}$ $u_{L}^{+}\exp\left[-\frac{i}{\hbar}\left(\varepsilon_{p}t-pz\right)\right],$
$\psi_{R}^{-}={}$ $u_{R}^{-}\exp\left[\frac{i}{\hbar}\left(\varepsilon_{p}t-pz\right)\right],$ $\psi_{L}^{-}={}$ $u_{L}^{-}\exp\left[\frac{i}{\hbar}\left(\varepsilon_{p}t-pz\right)\right].$ (100)
@@
$\Lambda=\mathbf{S}\cdot\frac{\boldsymbol{p}}{\left|\boldsymbol{p}\right|}=\frac{\hbar}{2}\left(\mathbf{\boldsymbol{\Sigma}}\cdot\frac{\boldsymbol{p}}{\left|\boldsymbol{p}\right|}\right),$ (103)
where the inner product represents the projection of spin onto the momentum. Therefore, when $\mathbf{S}$ and $\boldsymbol{p}$ are orthogonal, their inner product is zero.
@@
$\Lambda=\mathbf{S}\cdot\frac{\boldsymbol{p}}{\left|\boldsymbol{p}\right|}=\frac{1}{\left|\boldsymbol{p}\right|}S_{z}=\frac{\hbar}{2\left|\boldsymbol{p}\right|}\Sigma{}_{z},$ (104)
that is, the projection of $\mathbf{S}$ onto $\boldsymbol{p}$ is precisely the $S_{z}$ component of the spin. We have for $\Sigma{}_{z}$, according to Eq. (102),
@@
$\Lambda u_{L}^{+}={}$ $\frac{\hbar}{2\left|\boldsymbol{p}\right|}N\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}\begin{pmatrix}0\\ 1\\ 0\\ -\frac{pc}{\varepsilon_{p}+mc^{2}}\end{pmatrix},$
[…]
@@
Figure 6: Illustrated depiction of helicity. In the left panel a) the alignment of momentum and spin in the same direction resulting in positive helicity, or right-handedness. Conversely, in the right panel b) the opposing directions of momentum and spin lead to negative helicity, or left-handedness. Adapted from Ref. [18].
Given that positive energy pertains to particles and negative energy pertains to antiparticles, we may put aside brevity in favor of a more explanatory notation of the solutions for a free particle. In the following, we will adopt the notation found in (VI), applying our findings collectively to the states.
@@
$-i\hbar\psi\frac{\partial\psi^{\dagger}}{\partial t}$ $=ic\hbar\psi\boldsymbol{\alpha}\cdot\mathbf{\nabla\psi^{\dagger}}+mc^{2}\psi\beta\psi^{\dagger}.$ (109)
@@
that we can use to calculate the normalization factor $N$ of the free particle solutions. Taking the bispinor in Eq. (VI) associated with $u_{R}^{+}$ as an example, we compute

### v2

@@ IV Klein-Gordon equation
During 1926, a possible relativistic wave equation was derived by at least seven authors: Klein [50], Schrödinger [35], Fock [51], Donder, Dungen [52], Kudar [53] and Gordon [54]. Priority in the nomenclature is given to Oscar Klein, who proposed the equation in April 1926, while Fock provided a more intricate exploration of relativistic wave mechanics [3]. We will follow the more direct derivation, starting from the relativistic relation between energy $E$ and linear momentum $\mathbf{p}=(p_{1},p_{2},p_{3})$, expressed as
$E^{2}=(\mathbf{p}c)^{2}+(mc^{2})^{2},$ (37)
where the potential energy $V(\mathbf{r},t)$ is taken as zero and $m$ refers to the rest mass. We can replace both $E$ and $\mathbf{p}$ quantities in Eq. (37) by operators the operator in Eqs. (6) and (7), so that we obtain
@@
$\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}\psi\left(\mathbf{r},t\right)=\boldsymbol{\nabla}^{2}\psi\left(\mathbf{r},t\right)-\left(\frac{mc}{\hbar}\right)^{2}\psi\left(\mathbf{r},t\right),$ (39)
@@
Another crucial aspect of the KG equation is investigating its applicability in physical systems, particularly considering the range of problems that can be addressed with the Schrödinger equation. Therefore, applying the KG equation to a simple system and testing its validity is only fair. For instance, as an elementary problem, let us analyze the energy spectrum of a free particle. Taking Eq. (40) and simplifying it, we have
$\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}\Psi\left(\mathbf{r},t\right)-\boldsymbol{\nabla}^{2}\Psi\left(\mathbf{r},t\right)+\mu{{}^{2}}\Psi\left(\mathbf{r},t\right)=0,$ (46)
which is expected to have plane wave solutions with a time dependency of the type $(i\varepsilon t)/\hbar$ and a spatial dependence of the type $(i\mathbf{p}\cdot\mathbf{r})/\hbar$, so that
$\Psi\left(\mathbf{r},t\right)=N\exp[-\frac{i}{\hbar}(\varepsilon t-\mathbf{p}\cdot\mathbf{r})],$ (47)
@@
$\varepsilon^{2}=\mathbf{p}^{2}c^{2}+\mu^{2}c^{2}\hbar^{2},$ (48)
@@
where we note that, although it is a preserved quantity, $\rho(\mathbf{r},t)$ is not positive definite, unlike Eq. (14). This posed an immediate challenge as it complicated the probabilistic interpretation of the wave function. Nevertheless, we now know that the correct interpretation is that $\rho(\mathbf{r},t)$ is a probability charge density [41].
@@ V Pauli equation
An important event in the search for a relativistic wave equation was the discovery of spin, initially believed to be intrinsically relativistic. The first indication of spin’s existence arose from the Stern-Gerlach (SG) experiment [58], where an atomic beam of silver was subjected to an inhomogeneous magnetic field. Silver’s 47th electron’s magnetic moment approximates the atom’s magnetic moment because of its specific structure. If the electron behaved classically, its magnetic moment would exhibit a continuous range of values, resulting in a Gaussian pattern after passing through the magnetic field. However, in the SG experiment, deviations were observed for all atoms, with detection occurring in two preferential regions, as illustrated in Fig. 4. This suggested that the magnetic moment of the electron had only two directions, a phenomenon which Sommerfeld interpreted as evidence of space quantization [23].
Figure 4: SG experiment illustrated schematically. The beam of silver atoms is expelled from the furnace, passes through a collimator, and is subjected to a heterogeneous magnetic field $\mathbf{B}$. In the detector, two preferential regions are perceived.
@@
First, let us discuss how the Schrödinger equation deals with electromagnetic interactions by writing it as a function of $\mathbf{p}$ as
@@
$\boldsymbol{\Pi}=\mathbf{p}-\frac{e}{c}\mathbf{A},$ (51)
@@
$\mathbf{p}$ $\rightarrow\boldsymbol{\Pi}=\mathbf{p}-\frac{e}{c}\mathbf{A},$ (52)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left[\frac{1}{2m}\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)^{2}+e\phi\right]\psi\left(\mathbf{r},t\right).$ (54)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left\{\frac{1}{2m}\left[\mathbf{\boldsymbol{\sigma}}\ \cdot\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)\right]^{2}+e\phi\right\}\psi\left(\mathbf{r},t\right),$ (55)
@@
understood as operators associated with spin. In this equation, $\psi\left(\mathbf{r},t\right)$ is a two-component wave function [62] – a mathematical entity called a spinor represented by a column matrix with two components. Incorporating the spinor into the wave function is a direct consequence of the binary nature of spin. Consequently, the solutions of the Pauli equation involve two differential equations, as opposed to one in the case of the Schrödinger equation. In this work, we will not go into the formal aspects of the spinor, which can be found in [63].
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left\{\frac{1}{2m}\left[\left(\mathbf{p}-\frac{e}{c}\mathbf{A}\right)^{2}-\frac{e\hbar}{c}\mathbf{\boldsymbol{\sigma}}\cdot\mathbf{B}\right]+e\phi\right\}\psi\left(\mathbf{r},t\right),$ (58)
where $\mathbf{B}$ is the magnetic field, defined as $\mathbf{B}=\boldsymbol{\nabla}\times\mathbf{A}$. In the context of our discussion, the significance of the Pauli equation must be underscored. By incorporating spin, even if forcefully, into the Schrödinger wave function, his model became the most accurate representation of an electron interacting with an electromagnetic field under a non-relativistic regime.
@@ VI Dirac equation
Dirac, dissatisfied with the KG equation, believed that a linear relativistic wave equation could be developed in terms of the temporal derivative. This conviction arose from transformation theory [15], a concept to which he and Jordan contributed, aimed at generalizing both matrix and wave mechanics [23]. Upon comparing the Schrödinger equation with the KG equation, it is apparent that the Schrödinger equation fulfills this criterion, whereas the KG equation does not. To understanding why this linearity is vital in the development of quantum theory, see [66, 67]. Dirac expressed his strong preference for transformation theory with his statement: “The transformation theory had become my darling. I was not interested in considering any theory that would not fit my darling” [68] as cited in [15]. Notably, since the Dirac equation is of utmost importance for this article, the steps concerning it will be elaborated upon in greater detail, with a deeper exploration of specific topics compared to the earlier models. The derivation presented here emphasizes Dirac’s initial insight, while a more didactic derivation can be found in [69] and more explicitly detailed in [12].
Thus, while “playing with the equations” [15], Dirac obtained
$\mathbf{\mathbf{p}}^{2}\cdot\mathbbm{1}_{2}=(\boldsymbol{\sigma}\cdot\mathbf{p})^{2}.$ (59)
@@
Dirac claimed that both he and Pauli derived those matrices independently [15].
@@
$\gamma_{\mu}\gamma_{\nu}={}$ $-\gamma_{\nu}\gamma_{\mu},$ $\mbox{for }\mu\neq\nu.$
@@
which we can rewrite using the scalar product between $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $\mathbf{p}$ as
$E=c\boldsymbol{\alpha}\cdot\mathbf{p}+\beta mc^{2},$ (74)
@@
$i\hbar\frac{\partial}{\partial t}=c\boldsymbol{\alpha}\cdot\mathbf{p}+\beta mc^{2}.$ (75)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left(c\boldsymbol{\alpha}\cdot\mathbf{p}+\beta mc^{2}\right)\psi\left(\mathbf{r},t\right).$ (76)
@@
$H=c\boldsymbol{\alpha}\cdot\mathbf{p}+\beta mc^{2},$ (78)
@@
$i\hbar\frac{\partial}{\partial t}\psi\left(\mathbf{r},t\right)=\left(-i\hbar c\boldsymbol{\alpha}\cdot\nabla+\beta mc{{}^{2}}\right)\psi(\mathbf{r},t).$ (79)
@@
Just as we did with the KG equation, let us analyze the case of free particles. Firstly, we consider a particle at rest. Although it is a restricted case, it will provide us with the necessary insight regarding the allowed energy values. We employ the Dirac equation with $\mathbf{p}=0$, so that
@@
The first two expressions align with the anticipated outcomes for a free particle scenario, as suggested by the negative sign in the exponent. However, as we will explore in the subsequent case, the two lower components deviate from the expected behavior and are associated with negative energy.
@@
After some confusion [15], Dirac managed to formulate a coherent interpretation of this problem using the Pauli exclusion principle [72]. He conjectured a sea of negative energy, referred to as Dirac Sea, in which all the negative-energy states are filled, thus preventing an electron transition to negative energy. According to Dirac’s idea, a “hole” in this sea could behave like a particle with positive charge. Initially, Dirac speculated that the particle associated with the "hole" was the proton. This belief stemmed from the prevailing notion at the time that there were only two elementary particles in nature. However, as pointed out by Oppenheimer [73], the presence of a proton in this sea of negative energy would lead to annihilation and the release of energy in the form of photons, which means, in a practical sense, the instability of matter. Weyl [74], employing the symmetry in the Maxwell and Dirac equations, pointed out that the mass of the hole would have to be the same as the electron. In response to these critiques, Dirac stated, in an article published in 1931 [75], that if there is a hole, it represents a new, experimentally unknown particle with the same mass of the electron but opposite charge – an “antielectron”, so to speak. Today, this particle has been experimentally confirmed and is known as the positron. Therefore, the new degree of freedom observed in the bispinor arises precisely from the antiparticle associated with the particle. The energy diagram in Fig. 5 illustrates the concept of the Dirac Sea.
@@
Figure 5: Dirac equation’s energy diagram.
@@
$\Psi_{R}^{\pm}$ $=u_{R}^{\pm}\exp\left[\mp\frac{i}{\hbar}\left(\varepsilon_{p}t-pz\right)\right],$ (101)
$\Psi_{L}^{\pm}$ $=u_{L}^{\pm}\exp\left[\mp\frac{i}{\hbar}\left(\varepsilon_{p}t-pz\right)\right].$
@@
$\Lambda=\mathbf{S}\cdot\frac{\mathbf{p}}{\left|\mathbf{p}\right|}=\frac{\hbar}{2}\left(\mathbf{\boldsymbol{\Sigma}}\cdot\frac{\mathbf{p}}{\left|\mathbf{p}\right|}\right),$ (104)
where the inner product represents the projection of spin onto the momentum. Therefore, when $\mathbf{S}$ and $\mathbf{p}$ are orthogonal, their inner product is zero.
@@
$\Lambda=\mathbf{S}\cdot\frac{\mathbf{p}}{\left|\mathbf{p}\right|}=\frac{1}{\left|\mathbf{p}\right|}S_{z}=\frac{\hbar}{2\left|\mathbf{p}\right|}\Sigma{}_{z}$, (105)
that is, the projection of $\mathbf{S}$ onto $\mathbf{p}$ is precisely the $S_{z}$ component of the spin. We have for $\Sigma{}_{z}$, according to Eq. (103),
@@
$\Lambda u_{L}^{+}={}$ $\frac{\hbar}{2\left|\mathbf{p}\right|}N\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}\begin{pmatrix}0\\ 1\\ 0\\ -\frac{pc}{\varepsilon_{p}+mc^{2}}\end{pmatrix}$,
$={}$ $-\frac{\hbar}{2\left|\mathbf{p}\right|}u_{L}^{+},$ (107)
where the negative sign denotes negative helicity, commonly known as left-handed helicity. Applying the helicity operator to a positive, or right-handed, helicity solution would result in the constant (${\hbar}/{2\left|\mathbf{p}\right|}$) being multiplied by a positive sign. The distinct sign and its associated helicity confer the spin attribute to the respective solution. It is worth noting that we employ $u_{L}^{+}$ for simplicity in this example, but the same helicity would hold for $\psi_{L}^{+}$ since what holds significance is the sign acquired during the measurement process. Thus, the exponential term in $\psi_{L}^{+}$ does not affect the outcome of the measurement.
@@
Figure 6: Illustrated depiction of helicity. In the left panel a) the alignment of momentum and spin in the same direction resulting in positive helicity, or right-handedness. Conversely, in the right panel b) the opposing directions of momentum and spin lead to negative helicity, or left-handedness.
Given that positive energy pertains to particles and negative energy pertains to antiparticles, we may put aside brevity in favor of a more explanatory notation of the solutions for a free particle. In the following, we will adopt the notation found in (101), applying our findings collectively to the states
@@
$i\hbar\psi\frac{\partial\psi^{\dagger}}{\partial t}=ic\hbar\psi\boldsymbol{\alpha}\cdot\mathbf{\nabla\psi^{\dagger}}+mc^{2}\psi\beta\psi^{\dagger}.$ (110)
@@
that we can use to calculate the normalization factor $N$ of the free particle solutions. Taking the bispinor in Eq. (101) associated with $u_{R}^{+}$ as an example, we compute
@@
We add that the modern attitude toward the Dirac equation differs from the one presented here, where we interpret it in a manner similar to the Schrödinger equation. This discrepancy arises because the “single-particle” interpretation breaks down, as the creation and annihilation of particles become possible in nontrivial interactions. The proper framework to address this issue is quantum field theory (QFT) [41, 13].